Locate the absolute maximum and minimum for each of the following functions and justify your responses. (Show all work and justify)


 𝑓(π‘₯) = (Cube root of x^2) βˆ’ π‘₯ on the interval [βˆ’1,1]


1
Expert's answer
2021-02-05T04:57:01-0500

Given: "f(x)=\sqrt[3]{x^2}-x" , "\left [ -1,1 \right ]"

Find the first derivative of the given function.

"f(x)=\sqrt[3]{x^2}-x \Rightarrow f(x) =x^{\frac{2}{3}}-x"

"\Rightarrow f'(x)=\frac{2}{3}x^{\frac{-1}{3}}-1"

For critical values, set "f'(x)=0\Rightarrow \frac{2}{3}x^{\frac{-1}{3}}-1=0"

"\Rightarrow x^{\frac{-1}{3}}=\frac{3}{2}\Rightarrow x=\left ( \frac{2}{3} \right )^3=\frac{8}{27}"

Also observe that "x=\frac{8}{27}\in \left [ -1,1 \right ]"

Now find the function values at the end values "x=-1,1"

and at the critical value "x=\frac{8}{27}"

"x=-1\Rightarrow f(-1)=(-1)^{\frac{2}{3}}-(-1)=1+1=2" (Absolute maximum)

"x=1\Rightarrow f(1)=(1)^{\frac{2}{3}}-(1)=1-1=0" (Absolute minimum)

"x=\frac{8}{27}\Rightarrow f(\frac{8}{27})=(\frac{8}{27})^{\frac{2}{3}}-(1)=\frac{4}{9}"

Therefore "f(x)" has absolute maximum at "x=-1" and

absolute maximum value "=2"

"f(x)" has absolute minimum value at "x=1" and

absolute minimum value "=0"